Ask
your own question, for FREE!
Mathematics
12 Online
The probability that a randomly chosen student at the University of New harmony is a woman is 0.60. the probablity that the student is studying education is 0.20. The conditional probability that the student is a woman, given that the student is studyingeducation, is 0.8. What is the conditional probability that the student is studying education, given that she is a woman? Coul anyone anwer this with the calculations included?
Still Need Help?
Join the QuestionCove community and study together with friends!
Does no one has any knowledge about statistics? XD i know the subject sucks but someone needs to know it right? :P I give a medal!!
P(W) = 0.6 P(E) = 0.2 P(W|E) = 0.8 |dw:1414963732563:dw| P(E|W) = 0.2 * 0.8 = 0.16
|dw:1414964003958:dw|
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
10 hours ago
0 Replies
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
1 day ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal